Find the slope of a line through the points (2, 5) and (3, 5).
1
Expert's answer
2017-03-22T08:35:05-0400
Find the slope of a line through the points (2,5) and (3,5).
We know that line y = kx + b is represented in the slope-intercept form, where k - slope.
We can find k and b.
(2,5) belongs to the line so 5 = 2k + b.
(3,5) belongs to the line so 5 = 3k + b.
we have that 2k + b = 5 = 3k + b then 2k + b = 3k + b and finaly k = 0.
We get that the slope of the line is equal to 0.